Exercise 23.32

A very long insulating cylinder of charge of radius 2.40cm carries a uniform linear density of 18.0nC/m .

i tried 6900 cm and140cm 1.4cm and1.3cm

Part A

If you put one probe of a voltmeter at the surface, how far from the surface must the other probe be placed so that the voltmeter reads 195V ?
Answer #1

E = -dV/dr

dV = -E*dr

dV = -(2*k*lamda/r)*dr

do integration both sides

v2 - V1 = inegral -(2*k*lamda/r)*dr

= -2*k*lamda*ln(r)

-195 = -2*k*lamda*ln*(r2 - r1)

ln(r2 - r1) = 195/(2*10^9*18*10^-9)

ln(r2 - 0.024) = 5.417

r2 - 0.024 = e^5.417

r2 = 0.024 + e^(5.417)

= 225.5 m <<<<<<------------Answer